solve by completing the square
2x^2+5-1=0

The solution are/is x=

Answer by stanbon(75887)

solve by completing the square
2x^2+5x-1=0
2x^2 + 5x = 1
2(x^2 + (5/2)x + (5/4)^2) = 1 + 2(5/4)^2
2(x+(5/4))^2 = 1 + 2(25/16) = 1 + (25/8) = 33/8
(x+(5/4))^2 = 33/16
Take the square root:
x+(5/4) = [sqrt33]/4 or x+(5/4) = -[sqrt33]/4
x = [-5+sqrt(33)]/4 or x = [-5-sqrt(33)]/4
============================================
Cheers,
Stan H.